paula is currently 8 years younger than twice dawns age. in 4 years the differences between their ages will be 4 years. how old are they now?
Found 2 solutions by ankor@dixie-net.com, Alan3354:
Answer by ankor@dixie-net.com(22740)

paula is currently 8 years younger than twice dawns age.
p = 2d - 8
in 4 years the differences between their ages will be 4 years.
This is dumb, the difference in their ages will always be 4 years
p = d + 4
how old are they now?
Replace p with (2d-8) from the first equation
2d - 8 = d + 4
2d - d = 4 + 8
d = 12 yrs is Dawn's age
then
12 + 4 = 16 yrs is Paul's age

Answer by Alan3354(69443)

paula is currently 8 years younger
-----------
How does that differ from "...paula is 8 years younger...?
What does currently contribute?